What volume will 2.50 mol of hydrogen (H2) occupy at -20.0 °C and 1.5 atm?
1
Expert's answer
2020-10-09T13:46:12-0400

Use ideal Gas Law

PV=nRT

"V=\frac{nRT}{P}=\frac{2.50 mol\times 0.082\frac{L\times atm}{mol\times K}\times(-20+273)}{1.5 atm} = 34.6 dm^3"